A bank's loan officer rates applicants for credit approval. The ratings are normally distributed with a mean of 200 and a standard deviation of 50. If an applicant is randomly selected, find the probability of a rating between 225 and 275.
Answer by stanbon(75887)

A bank's loan officer rates applicants for credit approval. The ratings are normally distributed with a mean of 200 and a standard deviation of 50. If an applicant is randomly selected, find the probability of a rating between 225 and 275.
---
z(225) = (225-200)/50 = 1/2
z(275) = (275-200)/50 = 75/50 = 3/2
------
P(225 < x < 275) = P(1/2 < z < 3/2) = normalcdf(1/2,3/2) = 0.2417
---------
Cheers,
Stan H.
-----------